OpenStudy (anonymous):

Differentiate: y = (v^3 - 2v[sqrt[v]])/v

OpenStudy (badhi):

\[y =\frac{v^3-2v\sqrt{v}}{v} = v^2-2\sqrt v\] I guess its easier to differentiate like this